# FreeStyle: Free Lunch for Text-guided Style Transfer using Diffusion Models

### v2

@@ Abstract
The rapid development of generative diffusion models has significantly advanced the field of style transfer. However, most current style transfer methods based on diffusion models typically involve a slow iterative optimization process, e.g., model fine-tuning and textual inversion of style concept. In this paper, we introduce FreeStyle, an innovative style transfer method built upon a pre-trained large diffusion model, requiring no further optimization. Besides, our method enables style transfer only through a text description of the desired style, eliminating the necessity of style images. Specifically, we propose a dual-stream encoder and single-stream decoder architecture, replacing the conventional U-Net in diffusion models. In the dual-stream encoder, two distinct branches take the content image and style text prompt as inputs, achieving content and style decoupling. In the decoder, we further modulate features from the dual streams based on a given content image and the corresponding style text prompt for precise style transfer. Our experimental results demonstrate high-quality synthesis and fidelity of our method across various content images and style text prompts. Compared with state-of-the-art methods that require training, our FreeStyle approach notably reduces the computational burden by thousands of iterations, while achieving comparable or superior performance across multiple evaluation metrics including CLIP Aesthetic Score, CLIP Score, and Preference. We have released the code anonymously at: https://anonymous.4open.science/r/FreeStyleAnonymous-0F9B.
@@ 1 Introduction
Image style transfer [1, 2] intends to transfer the natural image into the desired artistic image while preserving the content information. With the recent rapid development of generative diffusion models [3, 4], image style transfer has also witnessed significant advancements. These methods can be broadly classified into two categories: finetuning-based methods [5, 6] and inversion-based methods [7, 8]. The former (depicted in Figure 1 (a)) requires optimizing some or all parameters to degrade the model to generate images of specific styles, while the latter (illustrated in Figure 1 (b)) involves learning the specific style concept as the textual token to guide style-specific generation. Both approaches often require thousands or even more iterations of training, leading to significant computational costs and a slow optimization process.
Large text-guided diffusion models [3, 9, 10], on the other hand, are typically trained on large-scale datasets of text-image pairs, e.g., LAION dataset [11], which encompasses various style images and corresponding style text prompts. Consequently, these models [3, 4] inherently possess the generative ability for specific styles. Recent works [12, 13] have introduced a cross-image attention mechanism to pre-trained diffusion models, enabling control of appearance or style transfer without optimization. However, the use of appearance images or style images as references is still required. In some applications, users may not have access to reference images but want to engage in image transfer based on style text prompts. For instance, users can envision transforming their photos into styles reminiscent of Picasso or Da Vinci without possessing works by these renowned artists.
In this paper, we present a novel style transfer approach that requires neither optimization nor style images. Specifically, we propose a novel structure composed of a dual-stream encoder and a single-stream decoder. In this configuration, the dual-stream encoder separately encodes the content image and style text prompt as inputs, extracting features from the corresponding modalities for integration in the decoder. It has been demonstrated that the low-frequency signals and high-frequency signals of an image are strongly correlated with its semantic information and style information [14, 15], respectively. We instantiate two modulation factors to balance low-frequency features from the U-Net’s main backbone and high-frequency features from skip connections to implement image style transfer. The first scaling factor regulates the strength of style transfer in the image and the second scaling factor controls the degree of content preservation in the images. Our approach is extremely simple and efficient, requiring only the adjustment of appropriate scaling factors to achieve the transfer of a specific style for any image.
@@ 2 Related Work
Image Style Transfer. The field of style transfer plays a pivotal role in image processing and computer vision. It has seen a rapid evolution from manual texture synthesis [21, 22] to advanced neural style transfer (NST) [1, 23, 24, 25, 19, 18], marking a significant shift from traditional techniques to modern deep learning approaches. Generative Adversarial Networks (GANs) [26], with impressive image generation capabilities, have been rapidly applied to style transfer tasks [27, 28, 29], further advancing the development of the field. With the recent rapid development of generative diffusion models [30, 31], significant progress has been made in image style transfer. These techniques can be classified into two main categories: finetune-based methods and inversion-based methods. Finetune-based methods [5, 6] optimize some or all of the model parameters using extensive style images, embedding their visual style into the model’s output domain. In contrast, inversion-based methods [7, 8] embed style or content concepts into special word embeddings using style or content images and achieve style transfer with prompts containing these word embeddings. The aforementioned methods based on diffusion models require style images for training models, resulting in a slow optimization process. Recent works [12, 13] introduce a cross-image attention mechanism and develop a style transfer method that does not require any optimization. However, these methods still rely on style images as references. As a text-guided style transfer method, FreeStyle differs by modulating features of the diffusion model, leveraging its inherent decoupling ability for style transformation without the need for extra optimization or style reference images.
@@
Text-guided Synthesis. GAN-CLS [32] is the first to achieve text-guided image synthesis of flowers and birds using recurrent neural networks [33] and Generative Adversarial Networks [26]. Subsequently, numerous efforts in text-guided image generation [34, 35, 36] have propelled rapid development in this field. Benefiting from the introduction of CLIP [37], the remarkable generative capabilities of text-to-image models [38, 10] have garnered significant attention from researchers, driven by the advancements in diffusion models. In addition to generating images that match text descriptions, text-guided techniques are now widely used in various tasks such as image editing [20, 39, 40], image restoration [41, 42], and video synthesis [43] etc. Tsu-Jui Fu et al. [44] argue that traditional style transfer methods, which depend on pre-prepared specific style images, are both inconvenient and creativity-limiting in practical applications. Following this, a new style transfer method that is guided by textual descriptions [17, 45] has been introduced, offering enhanced flexibility and convenience. This not only simplifies complex artistic creation but also makes advanced image manipulation accessible to a broader audience without the need for specialized graphic design skills. As a result, text-guided image processing is revolutionizing the way we interact with and create visual content.
@@
Deep Model Fusion. Deep model fusion [46, 47, 48] endeavors to integrate multiple deep neural networks (DNNs) into a singular network, maintaining their inherent capabilities and even surpassing the performance of multi-task training [49, 50]. With the emergence of new large language models (LLMs), such as GPT-3 [51], GPT-4 [52], T5 [53] and BERT [54], there is increasing attention on applying weighted averaging (WA) techniques [55, 56, 57] to these models. For instance, B-tuning [58] utilizes Bayesian learning to calculate posterior prediction distributions, thereby fine-tuning the top-K ranked pre-trained models based on their transferability. Zoo-tuning [59] aggregates the weights of pre-trained models with aligned channels to create a final model adapted to downstream tasks, addressing the high costs associated with migrating large models. For diffusion models, FreeU [60] strategically reweights the contributions of feature maps from U-Net’s skip connections and backbone to effectively enhance the quality of the generated images without any training. Surpassing FreeU, we further analyze and decouple the features of style text and content images. By directly integrating the two encoder branches of pre-trained U-Net in the diffusion model and modulating their features, we enable style image transfer without requiring textual input or additional training.
@@ 3.1 Preliminaries
Diffusion models [30, 31] involve a forward diffusion process and a reverse denoising process. During the forward process, Gaussian noise $\epsilon$ is progressively added to the clean sample $x_{0}$, with the intensity of the added noise $\epsilon$ increasing as $t\in\left[1,2,\dots,T\right]$ increases. The noised image at step $t$ is obtained through the diffusion process:
@@ 3.2 Model Structure of FreeStyle
In diffusion models, the U-Net structure is commonly used as the noise prediction network. It consists of an encoder and a decoder, along with skip connections that facilitate information exchange between corresponding layers of the encoder and decoder. Inspired by FreeU [60], which proposes to balance the low-frequency and high-frequency features from the U-Net’s backbone and skip layers, we introduce a novel modulate method of fusing content information and style information applied to style transfer. Figure 2 (a) illustrates the overall structure of FreeStyle, which consists of a dual-stream encoder and a single-stream decoder. The dual-stream encoder in FreeStyle comprises two U-Net encoders with shared parameters, while the single-stream decoder is made up of the U-Net decoder structure. The dual-stream downsampling process can be described as follows:
@@ 3.3 Feature Modulation Module
We strategically reweight the contributions of feature maps from the encoders of two parameter-sharing U-Nets, effectively leveraging the strengths of both components to implement image style transfer. It has been demonstrated that images consist of low-frequency signals controlling image content and high-frequency signals governing image style [14, 15]. We implement an effective training-free style transfer by modulating the style feature $f_{s}$ and the content feature $f_{c}$ to complete artistic image generation.
As shown in Figure 2(b), the content feature $f_{c}$ is generated guided by the noise-free content image $x_{0}$, while the style feature $f_{s}$ is generated guided by the style text prompt $p$ and the noise-added image $x_{\sigma}$. During the upsampling process in U-Net, the features $f_{c}$ primarily influence the semantic expression of the generated result, while the features $f_{s}$ have a greater impact on the high-frequency detail information of the result. Consequently, we engage in special modulation of $f_{s}$ and $f_{c}$ to further activate the intrinsic style reconstruction capability of U-Net. To enhance the semantic characteristics of the feature $f_{c}$, we amplify their variance. Specifically, we apply a weight parameter $b$ (where $b$>1) to certain channels of the feature to expand their variance, the process can be represent as:
@@ 4.1 Implementation Details
Figure 5: Qualitative comparison with several state-of-the-art image style transfer methods, e.g., CLIPstyler [17], UDT2I [20], CAST [18], and StyTr2 [19]. For the “SDXL*” approach, we initially add $\delta$ ($\delta$ is set to 850) steps of noise to the content image, then guide the denoising for $T$ steps using style prompts to generate a stylized image.
@@ 4.2 Experimental Result
Qualitative Results. To verify the robustness and generalization ability of FreeStyle, we conduct numerous style transfer experiments with various styles across different content. Figure 3 presents the style transfer effects of FreeStyle in the domains of buildings, landscapes, animals, etc. The experiments include style transfer in “Chinese Ink”, “Embroidery Art”, “Oil Painting”, “Watercolor Painting”, “Studio Ghibli”, “Cyberpunk”, “Pixel Punk”, “Wasteland” and “Sketching”. We showcase the results of applying style transfer to human portraits using FreeStyle, as in Figure 4 (left). In this figure, we also conduct style transfer experiments with multiple styles, including “Ufotable”, “Studio Ghibli”, “JOJO” and “Illumination Entertainment”. Observations indicate that FreeStyle is capable of providing accurate style information for the style transfer results while almost completely preserving the content information. For instance, the stylized results for “JOJO” maintain the structural information, while reasonably adjusting the image according to the character traits in the “JOJO” anime, like bold outlines, strong lines, and vibrant coloring. This achieves a more natural fusion and expression of both style and content. It is noteworthy that we perform style transfer on images using fine-grained styles from four animation categories in Figure 4 (left). Despite this, FreeStyle is still able to achieve style transfer results with high recognizability and accurate styling. Additionally, we apply multiple physical style transfers to various everyday items. As illustrated in Figure 4 (right), FreeStyle demonstrates excellent style transfer effects across these styles.
@@
Qualitative Comparisons. As shown in Figure 5, we conduct extensive comparative experiments with state-of-the-art methods, covering various styles and diverse content images. The results of “SDXL*” (e.g., rows 3,5) are highly sensitive to its hyperparameter $\delta$, resulting in significant variance in its stylized images. In contrast, FreeStyle achieves effective style transfer more consistently, demonstrating stronger robustness for images of varying content and across diverse style transfer tasks. The results show the apparent advantages of our method over others, as it can reasonably modify shapes (e.g., rows 1,2,6), brushstrokes (e.g., rows 1-5), lines (e.g., rows 3,4), and colors (e.g., rows 1-6) to achieve superior artistic effects. In comparisons between our method and several others, it is noticeable that our approach more accurately achieves style expression (e.g., rows 2,3,6), especially in styles that are more challenging to transfer. Results from CAST [18] and StyTr2 [19] are often marked by noticeable halo effects (e.g., rows 3,5,6) and are blurred (e.g., rows 2,6). In contrast, FreeStyle can produce clear stylized images without any noticeable halo effects. In summary, Figure 5 indicates that our method exhibits greater robustness, more accurate style expression, and more artistic style transfer effects.
@@
Quantitative Comparisons. To better evaluate our method, we employed multiple quantitative metrics for assessment, the results of which are presented in Table 1. For all comparison methods, we utilized their publicly available pretrained parameters for sampling. Following the widely used quantitative experimental setup [7, 17], we performed style transfers on 202 content images including landscapes, architecture, people, and animals, across 10 styles (“Chinese Ink”, “Illumination Entertainment”, “Embroidery Art”, “Graffiti Art”, “Impressionism”, “Oil Painting”, “Watercolor Painting”, “Cyberpunk”, “Studio Ghibli”, “Sketching”), resulting in a total of 2020 stylized images for each method. For the CLIP Score [37], we calculate the cosine similarity between the CLIP image embeddings and the prompt text embeddings. Using the prompt as a style description, we believe that a higher CLIP Score indicates a more accurate expression of style. The CLIP Aesthetic Score evaluates the quality, aesthetics, and artistic nature of images using a publicly available pre-trained art scoring model. A higher CLIP Aesthetic Score indicates that the fusion of style and content is more aesthetically pleasing. Training Cost refers to the product of the number of parameters that need to be optimized during the training phase and the recommended number of iterations in the corresponding method.
@@
[…]
FreeStyle (ours) 6.3148 25.615 0M 54.6%
Table 1: Quantitative comparisons with state-of-the-art methods are conducted, using CLIP Aesthetic Score, CLIP Score, Training Cost and human preferences as our evaluation criteria.
User Study. In addition to CLIP Aesthetic Score and CLIP score, we conduct a user study to evaluate the alignment of stylized images from six methods with human preferences. Forty participants from diverse fields and age groups are invited to evaluate 80 sets of different images, each transferred in various styles using 6 different methods. Participants are asked to select the best stylized image based on considerations of style expression, content preservation, and aesthetics. The “Preference” column in Table 1 reflects the percentage of instances where FreeStyle is ranked first by evaluators, showing its clear superiority over other methods.

### v3

@@ Abstract
The rapid development of generative diffusion models has significantly advanced the field of style transfer. However, most current style transfer methods based on diffusion models typically involve a slow iterative optimization process, e.g., model fine-tuning and textual inversion of style concept. In this paper, we introduce FreeStyle, an innovative style transfer method built upon a pre-trained large diffusion model, requiring no further optimization. Besides, our method enables style transfer only through a text description of the desired style, eliminating the necessity of style images. Specifically, we propose a dual-stream encoder and single-stream decoder architecture, replacing the conventional U-Net in diffusion models. In the dual-stream encoder, two distinct branches take the content image and style text prompt as inputs, achieving content and style decoupling. In the decoder, we further modulate features from the dual streams based on a given content image and the corresponding style text prompt for precise style transfer. Our experimental results demonstrate high-quality synthesis and fidelity of our method across various content images and style text prompts. Compared with state-of-the-art methods that require training, our FreeStyle approach notably reduces the computational burden by thousands of iterations, while achieving comparable or superior performance across multiple evaluation metrics including CLIP Aesthetic Score, CLIP Score, and Preference. We have released the code at: https://github.com/FreeStyleFreeLunch/FreeStyle.
keywords:
Generate model , Diffusion models , Style transfer , Training-free , U-Net
journal: Pattern Recognition
[…]
[label1]organization=School of Cyberspace Security at Sun Yat-sen University, city=Guangdong, postcode=518107, country=China
[…]
[label2]organization=Institute of Software,Chinese Academy of Sciences, city=Beijing, postcode=100190, country=China
[…]
[label3]organization=Information Technology Service Center of People’s Court, city=Beijing, postcode=100745, country=China
[…]
[label4]organization=School of Computer Science and Technology, Harbin Institute of Technology, city=Shandong, postcode=264209, country=China
{graphicalabstract}
{highlights}
A novel style transfer method (FreeStyle) based on diffusion models is proposed, leveraging pre-trained diffusion model parameters to achieve exceptional style transfer performance. This approach eliminates the need for style fine-tuning or inversion, significantly reducing computational costs.
The U-Net architecture within the diffusion model is explored, successfully decoupling image style and content. A dual-stream encoder is employed to separately encode style and content features, while a single-stream decoder merges these features multiple times during the upsampling process, achieving training-free transfer.
A feature modulation module is introduced, which scales and truncates content and style features in both the time and frequency domains, enabling precise control over the intensity of style and content representation. This provides flexibility in adjusting the strength of the transfer process.
@@ 1 Introduction
Image style transfer intends to transfer the natural image into the desired artistic image while preserving the content information. With the recent rapid development of generative diffusion models [1, 2, 3], image style transfer has also witnessed significant advancements. These methods can be broadly classified into two categories: finetuning-based methods [4] and inversion-based methods [5, 6]. The former (depicted in Fig. 1 (a)) requires optimizing some or all parameters to degrade the model to generate images of specific styles, while the latter (illustrated in Fig. 1 (b)) involves learning the specific style concept as the textual token to guide style-specific generation. Both approaches often require thousands or even more iterations of training, leading to significant computational costs and a slow optimization process.
Large text-guided diffusion models [1], on the other hand, are typically trained on large-scale datasets of text-image pairs, e.g., LAION dataset [7], which encompasses various style images and corresponding style text prompts. Consequently, these models [1, 2] inherently possess the generative ability for specific styles. Recent works [8] have introduced a cross-image attention mechanism to pre-trained diffusion models, enabling control of appearance or style transfer without optimization. However, the use of appearance images or style images as references is still required. In some applications, users may not have access to reference images but want to engage in image transfer based on style text prompts. For instance, users can envision transforming their photos into styles reminiscent of Picasso or Da Vinci without possessing works by these renowned artists.
In this paper, we present a novel style transfer approach that requires neither optimization nor style images. Specifically, we propose a novel structure composed of a dual-stream encoder and a single-stream decoder. In this configuration, the dual-stream encoder separately encodes the content image and style text prompt as inputs, extracting features from the corresponding modalities for integration in the decoder. It has been demonstrated that the low-frequency signals and high-frequency signals of an image are strongly correlated with its semantic information and style information [9], respectively. We instantiate two modulation factors to balance low-frequency features from the U-Net’s main backbone and high-frequency features from skip connections to implement image style transfer. The first scaling factor regulates the strength of style transfer in the image and the second scaling factor controls the degree of content preservation in the images. Our approach is extremely simple and efficient, requiring only the adjustment of appropriate scaling factors to achieve the transfer of a specific style for any image.
@@ 2.1 Image Style Transfer
The field of style transfer plays a pivotal role in image processing and computer vision. It has seen a rapid evolution from manual texture synthesis to advanced neural style transfer (NST) [15, 13, 12], marking a significant shift from traditional techniques to modern deep learning approaches. Generative Adversarial Networks (GANs) [16], with impressive image generation capabilities, have been rapidly applied to style transfer tasks [17], further advancing the development of the field. With the recent rapid development of generative diffusion models [18], significant progress has been made in image style transfer. These techniques can be classified into two main categories: finetune-based methods and inversion-based methods. Finetune-based methods [4] optimize some or all of the model parameters using extensive style images, embedding their visual style into the model’s output domain. In contrast, inversion-based methods [5, 6] embed style or content concepts into special word embeddings using style or content images and achieve style transfer with prompts containing these word embeddings. The aforementioned methods based on diffusion models require style images for training models, resulting in a slow optimization process. Recent works [8] introduce a cross-image attention mechanism and develop a style transfer method that does not require any optimization. However, these methods still rely on style images as references. As a text-guided style transfer method, FreeStyle differs by modulating features of the diffusion model, leveraging its inherent decoupling ability for style transformation without the need for extra optimization or style reference images.
2.2 Text-guided Synthesis
@@
GAN-CLS [19] is the first to achieve text-guided image synthesis of flowers and birds using recurrent neural networks [20] and Generative Adversarial Networks [16]. Subsequently, numerous efforts in text-guided image generation [21] have propelled rapid development in this field. Benefiting from the introduction of CLIP [22], the remarkable generative capabilities of text-to-image models [23, 24] have garnered significant attention from researchers, driven by the advancements in diffusion models. In addition to generating images that match text descriptions, text-guided techniques are now widely used in various tasks such as image editing [14, 25], image restoration [26], and video synthesis [27] etc. Tsu-Jui Fu et al. [28] argue that traditional style transfer methods, which depend on pre-prepared specific style images, are both inconvenient and creativity-limiting in practical applications. Following this, a new style transfer method that is guided by textual descriptions [11] has been introduced, offering enhanced flexibility and convenience. This not only simplifies complex artistic creation but also makes advanced image manipulation accessible to a broader audience without the need for specialized graphic design skills. As a result, text-guided image processing is revolutionizing the way we interact with and create visual content.
2.3 Deep Model Fusion
@@
Deep model fusion [29] endeavors to integrate multiple deep neural networks (DNNs) into a singular network, maintaining their inherent capabilities and even surpassing the performance of multi-task training [30]. With the emergence of new large language models (LLMs), such as GPT-3 [31], GPT-4 [32], T5 [33] and BERT [34], there is increasing attention on applying weighted averaging (WA) techniques [35] to these models. For instance, B-tuning [36] utilizes Bayesian learning to calculate posterior prediction distributions, thereby fine-tuning the top-K ranked pre-trained models based on their transferability. Zoo-tuning [37] aggregates the weights of pre-trained models with aligned channels to create a final model adapted to downstream tasks, addressing the high costs associated with migrating large models. For diffusion models, FreeU [38] strategically reweights the contributions of feature maps from U-Net’s skip connections and backbone to effectively enhance the quality of the generated images without any training. In FreeStyle, we fuse two latent space embeddings from different modality inputs and decode the latent space representation, which has absorbed information from both inputs, to generate an image that integrates both style and content information.
@@ 3.1 Preliminaries
Diffusion models [18] involve a forward diffusion process and a reverse denoising process. During the forward process, Gaussian noise $\epsilon$ is progressively added to the clean sample $x_{0}$, with the intensity of the added noise $\epsilon$ increasing as $t\in\left[1,2,\dots,T\right]$ increases. The noised image at step $t$ is obtained through the diffusion process:
@@ 3.2 Model Structure of FreeStyle
In diffusion models, the U-Net structure is commonly used as the noise prediction network. It consists of an encoder and a decoder, along with skip connections that facilitate information exchange between corresponding layers of the encoder and decoder. We propose a novel modulation method for fusing content and style information in style transfer by balancing the low-frequency and high-frequency features from the U-Net’s backbone and skip layers. Fig. 2 (a) illustrates the overall structure of FreeStyle, which consists of a dual-stream encoder and a single-stream decoder. The dual-stream encoder in FreeStyle comprises two U-Net encoders with shared parameters, while the single-stream decoder is made up of the U-Net decoder structure. The dual-stream downsampling process can be described as follows:
@@ 3.3 Feature Modulation Module
We strategically reweight the contributions of feature maps from the encoders of two parameter-sharing U-Nets, effectively leveraging the strengths of both components to implement image style transfer. It has been demonstrated that images consist of low-frequency signals controlling image content and high-frequency signals governing image style [9]. We implement an effective training-free style transfer by modulating the style feature $f_{s}$ and the content feature $f_{c}$ to complete artistic image generation.
As shown in Fig. 2(b), the content feature $f_{c}$ is generated guided by the noise-free content image $x_{0}$, while the style feature $f_{s}$ is generated guided by the style text prompt $p$ and the noise-added image $x_{\sigma}$. During the upsampling process in U-Net, the features $f_{c}$ primarily influence the semantic expression of the generated result, while the features $f_{s}$ have a greater impact on the high-frequency detail information of the result. Consequently, we engage in special modulation of $f_{s}$ and $f_{c}$ to further activate the intrinsic style reconstruction capability of U-Net. To enhance the semantic characteristics of the feature $f_{c}$, we amplify their variance. Specifically, we apply a weight parameter $b$ (where $b$ >1) to certain channels of the feature to expand their variance, the process can be represent as:
@@ 4.1 Implementation Details
Figure 5: Qualitative comparison with several state-of-the-art image style transfer methods, e.g., InST [5], CLIPstyler [11], UDT2I [14], CAST [12], StyTr2 [13] and DreamStyler [39].
@@ 4.2 Experimental Result
Qualitative Results. To verify the robustness and generalization ability of FreeStyle, we conduct numerous style transfer experiments with various styles across different content. Fig. 3 presents the style transfer effects of FreeStyle in the domains of buildings, landscapes, animals, etc. The experiments include style transfer in “Chinese Ink”, “Embroidery Art”, “Oil Painting”, “Watercolor Painting”, “Studio Ghibli”, “Cyberpunk”, “Pixel Punk”, “Wasteland” and “Sketching”. We showcase the results of applying style transfer to human portraits using FreeStyle, as in Fig. 4 (left). In this figure, we also conduct style transfer experiments with multiple styles, including “Ufotable”, “Studio Ghibli”, “JOJO” and “Illumination Entertainment”. Observations indicate that FreeStyle is capable of providing accurate style information for the style transfer results while almost completely preserving the content information. For instance, the stylized results for “JOJO” maintain the structural information, while reasonably adjusting the image according to the character traits in the “JOJO” anime, like bold outlines, strong lines, and vibrant coloring. This achieves a more natural fusion and expression of both style and content. It is noteworthy that we perform style transfer on images using fine-grained styles from four animation categories in Fig. 4 (left). Despite this, FreeStyle is still able to achieve style transfer results with high recognizability and accurate styling. Additionally, we apply multiple physical style transfers to various everyday items. As illustrated in Fig. 4 (right), FreeStyle demonstrates excellent style transfer effects across these styles.
@@
Qualitative Comparisons. As shown in Fig. 5, we conduct extensive comparative experiments with state-of-the-art methods, covering various styles and diverse content images. The results show the apparent advantages of our method over others, as it can reasonably modify shapes (e.g., rows 1,2,6), brushstrokes (e.g., rows 1-5), lines (e.g., rows 3,4), and colors (e.g., rows 1-6) to achieve superior artistic effects. In comparisons between our method and several others, it is noticeable that our approach more accurately achieves style expression (e.g., rows 2,3,6), especially in styles that are more challenging to transfer. In the results of the 5th line for both InST [5] and DreamStyler [39], varying degrees of leakage issues were observed. In contrast, FreeStyle avoids such problems by not using style images for the injection of style information. Additionally, compared to our method, style transfer results of InST [5] excessively and unnecessarily alter the content information. A key objective of style transfer tasks is to adapt to the target style while preserving the integrity of the content information as much as possible. Results from CAST [12] and StyTr2 [13] are often marked by noticeable halo effects (e.g., rows 3,5,6) and are blurred (e.g., rows 2,6). In contrast, FreeStyle can produce clear stylized images without any noticeable halo effects. The transfer results of both CLIPstyler [11] and UDT2I [14] exhibit issues of failed and inaccurate style expression. In summary, Fig. 5 indicates that our method exhibits greater robustness, more accurate style expression, and more artistic style transfer effects.
@@
Quantitative Comparisons. To better evaluate our method, we employed multiple quantitative metrics for assessment, the results of which are presented in Tab. 1. For all comparison methods, we utilized their publicly available pretrained parameters for sampling. Following the widely used quantitative experimental setup [5, 11], we performed style transfers on 202 content images including landscapes, architecture, people, and animals, across 10 styles (“Chinese Ink”, “Illumination Entertainment”, “Embroidery Art”, “Graffiti Art”, “Impressionism”, “Oil Painting”, “Watercolor Painting”, “Cyberpunk”, “Studio Ghibli”, “Sketching”), resulting in a total of 2020 stylized images for each method. For the CLIP Score [22], we calculate the cosine similarity between the CLIP image embeddings and the prompt text embeddings. Using the prompt as a style description, we believe that a higher CLIP Score indicates a more accurate expression of style. The CLIP Aesthetic Score evaluates the quality, aesthetics, and artistic nature of images using a publicly available pre-trained art scoring model. A higher CLIP Aesthetic Score indicates that the fusion of style and content is more aesthetically pleasing. Training Cost refers to the product of the number of parameters that need to be optimized during the training phase and the recommended number of iterations in the corresponding method. FreeStyle achieved state-of-the-art (SOTA) results in both the CLIP Aesthetic Score and Training Cost, as shown in Tab. 1. Additionally, FreeStyle demonstrated competitive results in the CLIP Score.
Table 1: Quantitative comparisons with state-of-the-art methods are conducted, using CLIP Aesthetic Score, CLIP Score and Training Cost as our evaluation criteria.
@@
CAST [12] 5.1462 22.347 3.51M×$\times$400
StyTr2 [13] 5.8613 22.300 35.39M×$\times$0.16M
CLIPstyler [11] 6.0275 27.614 0.62M×$\times$200
UDT2I [14] 6.2290 21.708 50$\times$ 10
FreeStyle (ours) 6.3148 25.615 0M